### v2

# Trustworthy AIGC Copyright Management with Full Lifecycle Recording and Multi-party Supervision in Blockchain

Abstract
As artificial intelligence technology becomes increasingly widespread, AI-generated content (AIGC) is gradually penetrating into many fields. Although AIGC plays an increasingly prominent role in business and cultural communication, the issue of copyright has also triggered widespread social discussion. The current legal system for copyright is built around human creators, yet in the realm of AIGC, the role of humans in content creation has diminished, with the creative expression primarily reliant on artificial intelligence. This discrepancy has led to numerous complexities and challenges in determining the copyright ownership of AIGC within the established legal boundaries. In view of this, it is necessary to meticulously record contributions of all entities involved in the generation of AIGC to achieve a fair distribution of copyright. For this purpose, this study thoroughly records the intermediate data generated throughout the full lifecycle of AIGC and deposits them into a decentralized blockchain system for secure multi-party supervision, thereby constructing a trustworthy AIGC copyright management system. In the event of copyright disputes, auditors can retrieve valuable proof from the blockchain, accurately defining the copyright ownership of AIGC products. Both theoretical and experimental analyses confirm that this scheme shows exceptional performance and security in the management of AIGC copyrights.
Index Terms:
AI-generated content (AIGC), Copyright management, Blockchain, Bloom filter, Blockchain-enabled auditing
1 Introduction
The widespread adoption of technologies such as AI painting and intelligent chat software like ChatGPT [^1] https://openai.com/chatgpt/ has drawn increased attention to Artificial Intelligence-Generated Content (AIGC) and has led to its broad application across various domains[32, 36, 37]. AIGC leverages artificial intelligence techniques, including Generative Adversarial Networks (GAN) [30] and diffusion models, to train on existing datasets and generate new content based on specific input prompts. In 2018, an artwork created by artificial intelligence was sold at Christie’s auction house for \$432,500222https://edition.cnn.com/style/article/obvious-ai-art-christies-auction-smart-creativity/index.html, marking the first-ever sale of AIGC product worldwide. The price achieved underscored the commercial and cultural significance of AIGC. However, this auction sparked a significant debate on AIGC copyright ownership, involving the distribution of rights among model providers, users, and other possible creative contributors of AIGC [18].
Copyright refers to the legal rights of creators to their literary, artistic and scientific works, with the aim of protecting the creative intellectual output of creators from unauthorized use by others, and encouraging innovation and cultural development. Traditional copyright law assumes that works are created by human creators. while the creation of AIGC is the result of collaboration among multiple participants. For instance, model providers develop AI models through extensive data training and computation, while users contribute to the content generation process by supplying prompts based on their personal experience and needs, which are then used to operate the AI content generation through the model [38]. The role of humans in content creation has diminished, raising new questions about whether the copyright of AIGC belongs to the AI technology itself, model providers, users, or other relevant parties. This challenges the requirement of human author in traditional copyright law. Moreover, copyright law generally necessitates that a work demonstrates a degree of originality, yet determining whether AI-generated content satisfies this criterion and assessing the originality of such content pose complex and undefined challenges [19]. As a result, AIGC frequently encounters quandaries related to copyright infringement.
In November 2023, Beijing Internet Court concluded a copyright infringement lawsuit [^3] https://patentlyo.com/media/2023/12/Li-v-Liu-Beijing-Internet-Court-20231127-with-English-Translation.pdf, in which the plaintiff submitted a comprehensive record of prompts and parameters employed in the content creation process, effectively establishing that the generated images reflected human creative intellectual input. Consequently, the court ruled in favor of the plaintiff, granting them the ownership of copyright over the image. This case illustrates that it is necessary to accurately record the key parameters and operational steps involved in the generation of AIGC products to ensure the copyright ownership of the generated AIGC product. Such records can serve as critical evidence to substantiate the intellectual and knowledge-intensive efforts embedded in the generation process, thereby supporting the legitimacy of copyright claims.
However, traditional data storage methods, with their centralized architecture, are vulnerable to attacks and tampering [34]. It allows each participant in the process of AIGC to potentially submit altered or fabricated proof that favors their interests during copyright attribution, affecting the determination of copyright ownership for AIGC. To address this flaw, there is an urgent need for a more secure and transparent recording system that ensures the integrity of proof. In this context, the application of blockchain appears to be crucial. Blockchain, with its unique decentralized structure and advanced encryption algorithms, can provide AIGC with a highly robust, decentralized, and manageable end-to-end copyright management platform [27]. By uploading the details of AI-generated content creation, copyright declarations, and information on all critical steps to the blockchain, these records are secured against unauthorized alterations once recorded [4]. Moreover, the timestamp of blockchain assigns an exact temporal verification to each record, enhancing the integrity and credibility of the proof while also ensuring the immutability of the transaction timeline, thus providing strong technical support for the protection and legal defense of AIGC copyrights. The transparency and traceability of blockchain provide convenience for copyright regulation, making the oversight process more open and fair, significantly improving the trustworthiness of the entire AIGC ecosystem.
So far, significant advancements have been achieved in the application of blockchain for copyright protection [8, 9, 11, 12, 23, 29], particularly in enhancing the security and traceability of copyright information storage. However, the majority of these applications are tailored to scenarios where the attribution of copyrights is unambiguous. The copyright of each AIGC product involves the contributions of multiple participants throughout various stages of content generation, which are often difficult to delineate and quantify. Evidently, the existing blockchain-enabled copyright protection strategies have not adequately addressed this complexity. Currently, research efforts are underway to apply blockchain to the management of AIGC copyrights [7, 33]. Nevertheless, these studies typically operate on the assumption that the copyright of AIGC is owned by a single entity, with a primary focus on the copyright registration and transaction management of the product itself. These research endeavors have not yet comprehensively considered the contributions of various participants throughout the content generation and their implications on copyright throughout the full lifecycle of AIGC products.
Figure 1: Graphical Abstract of AIGC-Chain
In this paper, we facilitate trustworthy AIGC copyright management by meticulously recording intermediate data generated throughout the full lifecycle of AIGC and leveraging blockchain to establish a decentralized multi-party supervision framework. As illustrated in Fig. 1, we pre-recorded the intermediate data generated at each stage of the AIGC lifecycle on the AIGC-Chain. In the instance of copyright disputes, we draw upon the data from the decentralized ledger to provide proof, thus offering a solid foundation and support for resolving copyright conflicts. Specifically, we maintain a world state with eight attributes for each AIGC product, ensuring that intermediate data of the generation process is securely stored on the blockchain in an immutable manner. This methodology establishes a robust framework ensuring traceability, verifiability, and integrity of AIGC products, effectively overcoming the constraints inherent in conventional content generation paradigms. By moving beyond both the anthropocentric authorship model and reductive instrumental approaches, it creates a reliable infrastructure for synergistic human-AI co-creation. Our main contributions can be summarized as follows:
  • We introduce AIGC-Chain, a trustworthy copyright management framework for the full lifecycle of AIGC, meticulously designing five transaction phases required for documentation and details of their state transition on the blockchain, aiming to achieve holistic management and complete recording of the status of AIGC.
  • We establish a decentralized multi-party supervision framework with blockchain, which enables automated copyright proof collection. This framework is augmented by IPFS technology to safeguard the the security and reliable storage of AIGC copyright.
  • We develop CTrace, a copyright tracing method based on bloom filter, which greatly improves the efficiency and accuracy of transaction query and verification in blockchain, provides strong proof support for copyright disputes, and thus realizes trustworthy AIGC copyright management.
The remainder of this paper is organized as follows. Section 2 reviews existing works that are related to ours. Section 3 introduces the preliminary to be used in the proposed system. Section 4 defines the system model, threat model, and design goals of AIGC-Chain. Section 5 presents the detailed construction of AIGC-Chain. Section 6 assesses the achievement of design goals. Section 7 evaluates the performance of AIGC-Chain. Section 8 concludes this paper and points out the future research directions.
2 Related Work
This section aims to review and summarize the research on copyright management for AIGC and the application of blockchain technology in copyright management.
2.1 The Copyright of AIGC
AIGC encompasses novels, articles, images, music, and other forms of content that are produced by artificial intelligence technology[28]. In recent years, the popularity of AI in creating artistic works has surged to unprecedented levels. This rapid expansion has precipitated numerous intricate challenges at the confluence of human and machine creativity, attracting considerable focus within the realm of copyright law[41]. These developments have engendered several key issues as follows.
Copyright Ownership: Given that AIGC is produced by AI programs rather than human creators, its copyright has become a subject of debate. Some contend that the copyright of AIGC should be assigned to the developers or proprietors of the AI programs, whereas others argue that AIGC should be treated as a public resource, accessible for unrestricted use by anyone. GAFFAR et al.[19] discussed the issue of copyright ownership in AIGC from a legal perspective, and attempted to clarify the definition of copyright and the methods for determining its ownership in AIGC. Several countries, including the United Kingdom, Ireland, New Zealand, and India, have proactively initiated legal reforms to more effectively address copyright challenges related to works that incorporate computer-generated elements and those created with the assistance of artificial intelligence[20]. However, these statutes fail to thoroughly address situations where a computer generates work with little or no direct human input, which could result in a lacuna in the legal framework.
Copyright Infringement: The generation process of AIGC products may involve operations such as copying, altering, and transforming existing works, which may potentially lead to copyright infringement[21]. Wang et al.[26] meticulously summarized the security and privacy challenges inherent in AIGC, offering fresh insights for the research of AIGC. For instance, some images generated by AI may use existing images as training data without obtaining authorization from the original author, thereby triggering infringement disputes. Besides, attackers may leverage the assistance of bots to disseminate a large volume of information, republishing the work while pretending it is the original one[7]. Consumers are likely to first encounter the manipulated work provided by the attacker, thereby successfully altering the copyright ownership of the target work.
2.2 Blockchain-enabled Copyright Management
As a distributed ledger technology, blockchain has played a crucial role in the field of copyright protection, owing to its decentralized, tamper-proof, and transparent attributes[6]. Digital assets on the blockchain, such as NFTs, can be permanently recorded and verified, with their uniqueness and verifiability allowing them to flow between any network entities and generate value[5]. Iyengar et al. [3] performed a thorough examination to assess the advantages and difficulties of applying blockchain technology in supply chain management and associated fields. Savilev [24] identified several critical issues in copyright management, such as copyright transparency, control over digital copies, and licensing, and suggests that blockchain technology can effectively address these challenges.
Various methods for copyright protection leveraging blockchain technology have been proposed thus far[22, 8, 23, 11, 39, 40]. Wang et al. [12] had harnessed the strengths of zero-watermarking algorithms and blockchain to safeguard the copyright of images. Xiao et al.[10] proposed a fair payment scheme based on blockchain that has proven particularly effective in ensuring the secure exchange of copyrighted content. Chen et al.[9] introduced a blockchain-based proactive defense strategy that offers the benefits of streamlined copyright transfers and preemptive penalties for infringement. Additionally, Xiao et al.[22] have developed a blockchain-based media-sharing network equipped with asymmetric fingerprinting technology, which meticulously records the processes of media content sharing and trading, thereby facilitating secure and dependable media exchange. Nonetheless, the fundamental question in AIGC copyright discussions is about the ownership of AIGC products. In contrast, the copyrights of products mentioned above are clear, serving to protect the rights of the identified copyright owners.
Currently, research on blockchain-enabled copyright protection technologies for AIGC is still limited. Chen et al. [25] discussed how blockchain provides a secure and reliable platform for AIGC. Nevertheless, he has not introduced a specific blockchain-based solution that effectively addresses the copyright concerns in AIGC. OpenAI published AIGC chain [^4] https://www.aigcchain.io/about in 2022, which allows users to contribute resources for training distributed AIGC models and acquiring rewards. However, the establishment of the entire ecosystem is still in its infancy, primarily aimed at providing crowdsourcing platforms for training AIGC models, without addressing issues related to AIGC copyright protection and commerce. Liu et al.[33] designed a digital assets trading method based on price game and blockchain. Liu et al. [7] leveraged blockchain to provide management services for AIGC products in cloud networks. However, these studies have not yet addressed how to manage the data generated throughout the entire lifecycle of an AIGC product, from its creation onwards.
Figure 2: The Structure of Consortium Blockchain Network
3 Priliminaries
In this section, we begin by exploring the blockchain technology utilized within AIGC-Chain. Then, we introduce the Indistinguishable Bloom Filter, which is used in our tracing method for transactions.
3.1 Consortium Blockchain
A consortium chain is a blockchain network controlled by a limited number of organizations, known as consortium members. These members agree on the rules and protocols governing the network and work together to validate transactions and maintain the integrity of the ledger[17]. Unlike public blockchains, where anyone can join and participate, consortium chains require explicit invitation and membership.
As illustrated in Fig. 2, the blockchain network is composed of orderer nodes and peer nodes. The orderer node is responsible for sorting transactions and distributing them to all peer nodes on relevant channels. Peer nodes, in turn, host both the ledger and smart contracts. The ledger maintains a systematically organized and tamper-proof record of all transactions. Additionally, the ledger incorporates a state database that tracks the current status, known as the world state. Every transaction generates a collection of key-value pairs that are utilized to update the world state.
The client sends a transaction proposal to one or more orderer nodes in the network. Orderer nodes pre-sort the transaction and validate its legitimacy. After pre-sorting, orderer nodes distribute the transaction to all relevant peer nodes on the channel. Upon receiving the transaction, peer nodes execute query or modification operations specified in the transaction and send the results back to orderer nodes. Orderer nodes collect the transaction results from all peer nodes and confirm the transaction. The transaction is then added to the blockchain of the channel and synchronized among all peer nodes in the network. The client application can query the transaction result to confirm that the transaction has been executed correctly.
3.2 Indistinguishable Bloom Filter
Indistinguishable Bloom filter (IBF) [16] obfuscates data by simulating future unknown queries. As an extension of the Bloom Filter (BFs) [14], IBF ensures efficient and accurate data retrieval while safeguarding data privacy. As illustrated in Fig. 3, IBF is constructed from an $m$-bit Twin Bloom Filter[15]. Within each twin, one cell is designated as chosen and the other as unchosen, with the unchosen cell potentially holding either 0 or 1. IBF encompasses three primary algorithms: IBF.Setup, IBF.Initi, IBF.Insert, and IBF.Check.
  • IBF.Setup$(1^{\eta},k,m)\to\left\{H_{key},H\right\}$: Construct $k$ pseudo-random hash functions $H_{key}=\left\{h_{1},h_{2},\cdots,h_{k+1}\right\}$ using the keyed-hash message authentication code (HMAC) and a hash function $H(x)=\mathbb{H}\%2$, where $\eta$ is the security parameter, $k$ is the number of hash functions, and $\mathbb{H}$ is a pseudorandom hash function.
  • IBF.Initi($m,H_{key},H,\gamma\to IBF$) Calculate $H(h_{k+1}(j)\oplus\gamma)$ to determine which cell in each twin is chosen, where $\gamma$ is a random number and $j\in\{0,m-1\}$.
  • IBF.Insert$(IBF,v,H_{key},H,\gamma)\to IBF$: For each value $v\in V$, set $IBF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1$ and $IBF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0$, where $i\in\left[1,k\right]$.
  • IBF.Check$(IBF,v^{{}^{\prime}},\gamma)\to 0/1$: Check whether $IBF\left[(h_{i}(v^{{}^{\prime}}))\right]\left[H(h_{k+1}(h_{i}(v^{{}^{\prime}}))\oplus\gamma\right]=1$ for the queried value $v^{{}^{\prime}}$. If the values of $k$ chosen cells are 1, $v^{{}^{\prime}}$ is included in IBF; otherwise, not included.
The correctness of IBF is established if, for all parameters $\eta$, $m$, $k$, $\gamma$, and any value $v$, following the execution of IBF.Setup, IBF.Init, IBF.Insert, and IBF.Check, the function IBF.Check returns a value of 1 when $v^{\prime}=v$, otherwise
$Pr[IBF.Check(IBF,v^{{}^{\prime}},\gamma)]=0]=1-\textnormal{negl}_{\textnormal{IBF}}(v,k,|V|),$
where the false positive rate of IBF is identical to that of traditional BF, i.e., neglIBF($m,k,|V|$) = neglBF($m,k,|V|$) $\approx(1-(1-\frac{1}{m})^{k|V|})^{k}\approx(1-e^{\frac{-k|V|}{m}})^{k}$.
Figure 3: The Technical Principle of Indistinguishable Bloom Filter
4 Models
This section begins by introducing the roles within AIGC-Chain and provides a concise overview of the system’s process design. It then proceeds to elaborate on the threat model and design goals of the system.
4.1 System Model
Fig. 4 depicts the four key roles within AIGC-Chain: User, Provider, Consumer, and Auditor. These roles interact with AIGC-Chain by submitting transactions to record copyright information on the blockchain, facilitating secure management of AIGC copyright details. Here are the detailed descriptions of the four roles:
Figure 4: System Model
  • User: User is the initiator of AIGC generation, existing as a light node on AIGC-Chain. It selects an appropriate model and inputs a series of parameters and keywords to trigger the generation. User is the initial owner of the AIGC product, with the right to use, share, and transfer its AIGC product.
  • Provider: Provider owns the trained AIGC model and exists as a full node on AIGC-Chain. It can provide AIGC generation service to User and charge a fee. Provider is responsible for promptly recording the usage and generation activities of models on AIGC-Chain.
  • Consumer: Consumer is the user who plans to purchase or utilize the AIGC product, existing as a light node on AIGC-Chain. It can get information about the AIGC product from the chain and propose a transaction to modify the ownership of the AIGC product.
  • Auditor: Auditor represents copyright certification institutions, courts, and other authoritative bodies on AIGC-Chain, existing as a full node. It gathers proof from AIGC-Chain to adjudicate copyright issues related to AIGC and ensure fair and just outcomes.
4.2 Threat Model
The threats considered in our scheme are as follows, which come from four entities, i.e., User, Provider, Consumer.
  • User, Provider, and Consumer may intentionally submit false or misleading proof to claim copyright ownership over the AIGC product.
  • User and Provider might furnish the Auditor with erroneous or fabricated transaction ID, which could lead to incorrect copyright assessments for AIGC products, thereby misleading the Auditor and preventing the tracing of the accurate copyright information related to the AIGC products.
  • During the copyright audit of the AIGC product, Auditors may collude with others, potentially leading to erroneous or unfair audit outcomes.
The threats beyond the scope mentioned above fall outside the scope of our primary treatment. These include activities such as eavesdropping on communication channels and attacks on blockchain node servers. These issues can be addressed by incorporating general solutions into our framework.
4.3 Design Goals
In response to the provided system model and threat model, the intended design goals are summarized as follows:
  • Trustworthy: Throughout each phase of generation, the intermediate data outputs are meticulously and accurately uploaded to the AIGC-Chain. By harnessing the inherent immutability of blockchain, we guarantee the trustworthy of the copyright proof retrieved, facilitating the management of copyright ownership.
  • Traceability: All data generated throughout the full lifecycle of an AIGC product can be permanently recorded and cryptographically verifiable, eliminating risks of tampering or unauthorized modifications to copyright-related information.
  • Fairness: Authoritative regulatory entities such as copyright certification institutions and courts are introduced as Auditor in AIGC-Chain. When copyright disputes arise, these nodes can retrieve proof from AIGC-Chain to ensure unbiased and fair judgments regarding the copyrights of AIGC products.
5 The Proposed System
In this section, we delve into the specifics of the proposed system. We begin by providing a comprehensive overview. Subsequently, we introduce a novel approach for tracing data within the blockchain. We conclude by elaborating on the operational workflow and the corresponding state transitions within the proposed system. For improved readability, a concise compilation of the key notations used in this section, including their definitions, is provided in Table I.
TABLE I: NOTATIONS AND DESCRIPTIONS
Notations Descriptions
$P$ the AIGC product
$P^{{}^{\prime}}$ the AIGC product to be verified
$Args$ the prompt for generating AIGC product
Product ID the key for an AIGC product in the world state
Model ID the unique identifier for the used model
Description Hash the hash of the input prompt Args
Product Hash the hash of the AIGC product $P$
User Sign the signature of User
Model Sign the signature of Provider
Owner Sign the signature of the owner of the AIGC product
BF the index of all transactions of the AIGC product
Status the current status of the AIGC product
5.1 Overview
By storing data generated throughout the AIGC lifecycle in a complete and unchangeable form on a distributed blockchain, we can ensure the trustworthy and traceability of the sources of AIGC products. This also guarantees that in the event of copyright disputes, data stored on AIGC-Chain can be used to determine the ownership of AIGC products, thereby protecting the security of AIGC copyrights. As depicted in Fig. 4, AIGC-Chain comprises 5 core procedures: Registration, Content Generation, Data Uploading, Copyright Trading, and Copyright Management.
AIGC-Chain is deployed on a private Ethereum network, which maintains its data in a decentralized blockchain ledger composed of blocks and the world state. Each block contains records of all preceding transactions that have been successfully executed, while the world state preserves the latest data state in the form of key-value pairs.
Within AIGC-Chain, a world state of 8 items is preserved in the ledger of AIGC-Chain for each AIGC product, as illustrated in Fig. 4. This world state undergoes continuous updates throughout the lifecycle of AIGC products, responsive to a multitude of operations. Among them, Model ID serves as a unique identifier for the training models employed. Given that participants of the blockchain have full access to all data on the chain, we employ a strategy to protect the privacy of AIGC products. Instead of storing the original content, we use a hash value Description Hash, derived from the input Args to update the world state. Model Sign and User Sign respectively denote the signatures of the model owner and the producer of the AIGC product. Once established, these signatures are permanent and cannot be modified. Similarly, Owner Sign is the signature of the owner of the AIGC product. Before the AIGC product is sold, the owner and the producer are identical. Following the transaction, this attribute will be updated. Status records the current state of the AIGC product, such as generating, generated, etc. Additionally, BF is employed to index all transaction information related to the AIGC product on the blockchain, facilitating the authenticity management of copyright. Details of BF will be discussed in Section 5.2.
Figure 5: The Data Structure of the Transaction in Fabric
5.2 Efficient Tracing Method Based on Indistinguishable Bloom Filter
As illustrated in Fig. 6, the generation of each AIGC product is inherently associated with at least three categories of on-chain transactions. In the event of a copyright dispute, Auditors could access detailed copyright proof through Product ID and require all participants to submit copyright proofs of their involvement in the generation process, enabling fine-grained copyright traceability. It should be noted that while the input parameters Args in transactions can validate the authenticity of transactions, the transparent nature of blockchain allows potential attackers to forge transactions with identical Args in attempts to manipulate copyright attribution of AIGC products. Chen et al. [9] proposed a fast query algorithm based on a singly-linked list, which constructs a local hash index structure to store the hash values of copyright information in a singly linked list. Compared with the traditional linear retrieval approach requiring full-node traversal, this algorithm enables copyright ownership verification without parsing the entire blockchain transaction data. However, this centralized storage system necessitates continuous maintenance of offline index replicas, leading to increased local storage costs. More critically, it also brings the risk of a single point of failure.
To address this vulnerability, we propose CTtrace, an efficient copyright traceable verification mechanism based on indistinguishable bloom filters. It constructs an $m$-bit twin array BF for each blockchain-recorded AIGC product with all associated transactions mapped. By implementing hash mapping of input parameter Args, timestamp CreateTime, and random value Nonce—as depicted in Fig. 5CTtrace ensures that even if attackers duplicate Args, they cannot reconstruct legitimate transaction records with identical timestamps and nonces.
Specifically, as depicted in Algorithm 1, CTrace comprises three stages: Setup, Insert, and Check. The smart contract automatically executes these stages upon triggering transactions for copyright storage and querying.
1) Setup: Upon the receipt of an AIGC generation request from User, a dedicated array BF and a set of hash functions $\left\{H_{key},H\right\}$ are generated for product P. These elements are subsequently embedded within the transaction and recorded in the world state on the blockchain. Access to this information is granted exclusively to Provider, Consumers, and Auditors affiliated with product P.
2) Insert: When a transaction is initiated to update the status of product P, its BF in the world state will be revised to:
$BF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1,$
$BF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0,$
where $v=Args||CreateTime||Nounce$ is retrieved from the proposed transaction Tx, and $i\in[1,k]$, $k$ represents the number of hash mapping functions.
3) Check: To confirm the authenticity of the transaction $Tx^{\prime}$ received from participants, check $v^{\prime}=Args^{\prime}||CreateTime^{\prime}||Nounce^{\prime}$ within $Tx^{\prime}$ as:
$BF\left[(h_{i}(v^{{}^{\prime}}))\right]\left[H(h_{k+1}(h_{i}(v^{{}^{\prime}}))\oplus\gamma\right]=1.$
If $v^{\prime}$ of $k$ selected hash functions for BF is found to be 1, it indicates that $Tx^{\prime}$ is linked to the generation of product P.
Algorithm 1 Details of CTrace
  Input: BF in the world state, transaction $Tx$ related to product P, and transaction $Tx^{\prime}$ to be verified
  Output: Whether $Tx^{\prime}$ is associated with product P
  Setup:
  Generate two $m$-bit sets as $BF$;
  $H_{key}=\left\{h_{1},h_{2},\cdots,h_{k+1}\right\}$;
  $H(x)=\mathbb{H}\%2$;
  Update $BF$ in the world state for product P;
  Insert:
  Propose a transaction $Tx$ related to P;
  $v=Args||CreateTime||Nounce$;
  for $i=1;i\leq k;i++$ do
     $BF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1$;
     $BF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0$;
  end for
  Update $BF$ in the world state for product P;
  Check:
  Retrieve $Args^{\prime}$ and $CreateTime^{\prime}$ from $Tx^{\prime}$;
  $v=Args^{\prime}||CreateTime^{\prime}||Nounce^{\prime}$;
  for $i=1;i\leq k;i++$ do
     if $BF\left[(h_{i}(v^{{}^{\prime}}))\right]\left[H(h_{k+1}(h_{i}(v^{{}^{\prime}}))\oplus\gamma\right]==0$ then
        return 0;
     end if
  end for
  return 1;
For instance, in the content generation based on diffusion models, the core workflow can be decomposed into two sequential phases: Forward Diffusion and Reverse Denoising, which collaboratively accomplish content generation through a multi-round iterative process. During the Forward Diffusion phase, key parameters are immortalized on the blockchain, encompassing noise schedule ${\beta_{1},\beta_{2},\ldots,\beta_{T}}$, the set of noisy images ${x_{1},x_{2},\ldots,x_{T}}$, and noise samples $\epsilon$ that are introduced. In the subsequent Reverse Denoising phase, blockchain records the set of denoised images ${x_{T},x_{T-1},\ldots,x_{0}}$ and the set of evolving parameter for the predicted noise $\epsilon_{\theta}(x_{t},t)$. Furthermore, considering the request transaction $Tx_{req}$ and multiple content exchange transactions $Tx_{exchange}$, we assume that the number of associated transactions per generated product is constrained by $n\leq 20$. To address the correctness verification requirements, when the false positive rate threshold is set to 0.1%, probability model derivation yields the optimal parameter combination $(k=10,m=288)$ satisfying:
$P_{\textit{fp}}\approx\left(1-e^{-kn/m}\right)^{k}\leq 0.001,$
This configuration guarantees that the system upholds the necessary probabilistic assurances, striking a balance between computational efficiency and the reliability of verification.
5.3 Lifecycle Management of AIGC Product
In the AIGC-Chain, we store all data generated throughout the AIGC creation process in the form of transactions on the blockchain.
We incorporate User Sign into the world state of AIGC products, establishing a direct connection between each product and the personal information of producers. This enhances the traceability of the AIGC products, allowing for the efficient monitoring of user identity information, including that of the producer. As a result, the measure effectively prevents the misuse of AIGC models and significantly reduces the production of fraudulent or low-quality AIGC content. To prevent duplicate transactions of product copyrights, we add Owner Sign in the world state of AIGC products, which ensures that each copyright transaction is permanently recorded on the blockchain. Consumers can review the transaction history of the product and modify Owner Sign upon completion of the deal. The transaction process is contingent upon the validation of the ownership signature; only when the signature is successfully verified can the transaction proceed.
To elucidate the full lifecycle recording process of AIGC-Chain with greater clarity, we illustrate the process with the generation of AIGC products via diffusion models as an example. Fig. 6 graphically depicts the transition of the world state of the product, with the modifications since the last transaction highlighted in red.
In Registration, CA registers roles participating in AIGC-Chain as lightweight or full nodes. Content Generation takes the prompt provided by User to the system and assigns unique Product ID as key for the aigproduct P. Instead of storing the actual prompt, it saves the hash of the prompt as Description Hash in the world state. After completing the content generation, Provider sends the hash value and other related generation information of P through Data Uploading and returns P to User. Consumer and User execute copyright transactions and ownership transfers for AIGC product through Copyright Trading. In the event of a copyright dispute, Auditors can invoke Copyright Management to retrieve the information related to P from AIGC-Chain. Following this, the auditors can proceed to analyze and evaluate the copyright of the product. The detailed introductions of the above five stages are as follows:
Figure 6: Lifecycle Management on the Blockchain
1) Registration: Users participating in AIGC-Chain are required to register on the blockchain network, with digital certificates $DC$ and public-private key pairs $(pk,sk)$ being issued by the Certificate Authority (CA) on the chain. Only users who have completed CA authentication are granted access to view and participate in blockchain activities.
Within this architecture, User and Consumer participate as lightweight nodes, predominantly hosted on personal computers and mobile devices. They can engage with blockchain interactions without being responsible for data storage. In contrast, Provider and Auditor, armed with more robust computing and storage power, serve as full nodes within the network, tasked with the responsibility of synchronizing the entire data on the blockchain.
Algorithm 2 Details of Content Generation
  Input: Product ID, prompt $Args$
  Output: transaction $Tx_{req}$
  User:
  $\sigma_{1}=sig_{sk_{User}}(Args)$;
  Send $\sigma_{1}$ and $Args$ to Provider;
  Provider:
  Verify $\sigma_{1}$ with $pk_{User}$;
  $\sigma_{2}=sig_{sk_{Provider}}(Args)$;
  Send $\sigma_{2}$ back to User;
  User:
  Description $\textit{Description ID}=H(Args)$;
  Verify the legality of $\sigma_{2}$ with $pk_{Provider}$;
  Owner $\textit{Owner Sign}=sig_{sk_{User}}(Description~{}Hash)$;
  Propose transaction $Tx_{req}$ to update the world state with $key=\textit{Product ID}$;
  Retrieve $Txid_{req}$ from transaction $Tx_{req}$ and store it locally or within a decentralized storage system.
2) Content Generation: When using an AIGC model to generate product P, User inputs the prompt $Args$ into the model and selects a unique Product ID as key for P. Next, User calls the smart contract to initiate a transaction $Tx_{req}$ to update the world state for P with key. As illustrated in Fig. 6, transaction $Tx_{req}$ updates the world state by updating the Description Hash, User Sign, Owner Sign, BF, and Status. Since the world state is public to all participants on the consortium chain, we use the hash of the prompt Description Hash in $Tx_{req}$ to prevent the prompt from being copied or maliciously used by an attacker. The specific process of this stage is as follows:
  • User computes $\sigma_{1}=sig_{sk_{User}}(Args)$ and sends $\sigma_{1}$ along with $Args$ to Provider for content generation.
  • Provider verifies $\sigma_{1}$ with $pk_{User}$ of User. The successful verification suggests that the prompt $Args$ aligns with the intent of the User and has not been altered during transmission.
  • Provider generates $\sigma_{2}=sig_{sk_{Provider}}(Args)$ and sends it back to User. This indicates that Provider acknowledges the legitimacy of the received $Args$, enabling further content generation operations.
  • User selects a random hash function $H$ and computes Description $\textit{Description ID}=H(Args)$.
  • User verifies the legality of $\sigma_{2}$; a successful validation implies the formation of consensus between User and Provider.
  • User generates signature Owner Sign with Description Hash.
  • A transaction $Tx_{req}$ is proposed by User to update the world state with $key=\textit{Product ID}$, where Status is used to indicate the current state of P, which is prepared.
  • User retrieves the transaction ID $Txid_{req}$ from transaction $Tx_{req}$ and may choose to store it either locally or within a decentralized storage system such as IPFS [^5] https://www.ipfs.tech, which facilitates the simplification of the copyright management process.
Algorithm 3 Details of Data Uploading
  Input: Product ID, prompt $Args$, and ID of the chosen model $Model~{}ID$
  Output: product P, transaction $Tx_{gen}$, and transaction $Tx_{upload}$
  Provider:
  Generate product P based on the prompt $Args$;
  $\sigma_{3}=sig_{sk_{Provider}}(P)$;
  Send $\sigma_{3}$ and generated product P back to User;
  User:
  Verify $\sigma_{3}$ with $pk_{Provider}$;
  $\sigma_{4}=sig_{sk_{User}}(P)$;
  Send $\sigma_{4}$ back to Provider;
  Provider:
  Verify the legality of $\sigma_{4}$ with $pk_{User}$;
  Model $\textit{Model Sign}=sig_{sk_{Provider}}(steps||seed||Created~{}Date)$
  Propose transaction $Tx_{gen}$ to update the world state with $key=\textit{Product ID}$;
  Retrieve $Txid_{gen}$ from transaction $Tx_{gen}$ and store it locally or within a decentralized storage system.
  User:
  Product $\textit{Product Hash}=H(P)$;
  Propose transaction $Tx_{upload}$ to update the world state with $key=\textit{Product ID}$;
  Retrieve $Txid_{upload}$ from transaction $Tx_{upload}$ and store it locally or within a decentralized storage system.
3) Data Uploading: Provider generates product P with the AIGC model based on the prompt $Args$ provided by User. Subsequently, Provider transmits P to User and initiates a transaction $Tx_{gen}$ to update additional information related to the generation process in the world state. Upon receiving product P, User initiates transaction $Tx_{upload}$ to add product P into the world state. Fig. 6 illustrates the alteration in the world state following the execution of transactions $Tx_{gen}$ and $Tx_{upload}$. In this process, the use of Product Hash in place of the original product P is aimed at preventing potential copyright infringement or misuse, while also ensuring the security of the data in AIGC-Chain.
  • Provider generates product P based on the prompt $Args$ and then computes signature $\sigma_{3}=sig_{sk_{Provider}}(P)$ for it.
  • Provider sends both $\sigma_{3}$ and the generated product P back to User for verification.
  • User authenticates the validity of $\sigma_{3}$; only upon successful verification will User accept product P and proceed to sign it as $\sigma_{4}=sig_{sk_{User}}(P)$, which will be sent back to Provider.
  • Provider verifies the legality of $\sigma_{4}$; a successful validation implies the formation of consensus between User and Provider.
  • Provider generates signature Model Sign with data related to the generation of the AIGC product, including steps, seed, and Created Date, which are data used and produced during the AIGC generation process.
  • Provider proposes a transaction $Tx_{gen}$ to update generation data related to P with $key=\textit{Product ID}$, where Status has shifted from prepared to generating.
  • User computes Product $\textit{Product Hash}=H(P)$ for product P.
  • A transaction $Tx_{upload}$ is proposed by User to update the world state, where Status has transitioned from generating to generated.
  • User and Provider retrieve transaction ID $Txid_{gen}$ and $Txid_{upload}$ from the blockchain respectively and can choose to store them either locally or within a decentralized storage system such as IPFS.
4) Copyright Trading: When Consumer plans to purchase the copyright of product P, it initially retrieves product details from the world state and verifies the validity of product P. Following successful validation, Consumer makes a payment to the owner and initiates a transaction $Tx_{exchange}$ on the blockchain to update Owner Sign and BF within the global state. The specific details are as follows:
  • Consumer solicits Product ID of aiming product P and the public key $pk_{owner}$ from the owner.
  • Consumer retrieves Owner Sign and Status from the world state with Product ID, validating the legality of Owner Sign with $pk_{owner}$. Simultaneously, it is necessary to ensure that Status is not prepared or generating.
  • A transaction $Tx_{exchange}$ is proposed by Consumer to update the world state.
5) Copyright Management: For a copyright dispute, Auditor can retrieve Product ID of the disputed product P by consulting the relevant world state on the AIGC-Chain. Provider and Provider are required to provide proof to Auditor to establish their connection with the disputed product. Auditor will then assess the ownership of the copyright based on the proof gathered.
  • User submits the proof $\varphi$ to Auditor, where $\varphi=(pk_{User},Txid_{req},Txid_{upload},prompt)$.
  • Provider delivers the proof $\phi$ to Auditor, where $\phi=(pk_{Provider},Txid_{gen},metadata)$, $metadata$ refers to some auxiliary information within the generation process of AIGC, including execution logs, intermediate results, etc.
  • Auditor utilizes $pk_{User}$ and $pk_{Provider}$ to verify User Sign and Model Sign retrieved from the world state of the AIGC-Chain for Product ID, to assess the correctness of the identities associated with the product-related roles.
  • Auditor employs CTrace.Check to verify the correctness of $Txid_{req}$, $Txid_{upload}$, and $Txid_{gen}$. An output of 1 indicates that the transaction is indeed related to the generation process of the AIGC product.
  • Auditor conducts an in-depth manual review of the originality, compliance, ownership, and value of the AIGC product based on the information provided by User and Provider, as well as the product information stored on AIGC-Chain.
6 Theoretically Evaluation
This section conducts a theoretical evaluation of the performance of the designed system about the design goals outlined in Section 4.
Theorem 1. The copyright-related data for AIGC products retrieved from the blockchain can be utilized to authenticate the copyright ownership of an AIGC product.
Proof: Blockchain is a linked series of blocks, with data dispersed across a network of nodes in a distributed manner. This distribution necessitates that any alteration to the data must be reflected across the majority of nodes, a process that is prohibitively challenging in practice. Additionally, the integration of a new block into the blockchain requires consensus from all nodes, which involves recomputing the necessary tasks and securing the validation and agreement of participants in the network. This process is both resource-intensive and infeasible for malicious purposes. Consequently, the data recorded on the blockchain is reliable and inviolable. This inherent security and immutability make the blockchain an enduring and trustworthy platform for recording the full lifecycle of AIGC products, providing reliable, immutable information for effective copyright management of AIGC products.
Theorem 2. It is impossible for anyone to attempt to alter the copyright-related data of an AIGC product by fabricating transactions on the blockchain.
Proof: CTrace establishes a clear association between AIGC product and transaction ID provided by User and Provider. Since forged $Txid^{\prime}$ cannot pass CTrace.check(), preventing them from being mapped within the blockchain, thereby confirming their irrelevance to the AIGC product. This mechanism effectively hinders attackers from posting fraudulent transactions on the blockchain to interfere with the accurate determination of copyright ownership for AIGC products.
Theorem 3. Auditors are equipped to make fair copyright ownership decisions by the decentralized and unchangeable properties of the blockchain.
Proof: Auditors responsible for copyright management are composed of copyright certification institutions, courts, and other authoritative bodies, which also serve as nodes within the AIGC-Chain network. They rely on dependable, immutable data from the blockchain. The decentralized nature of the blockchain necessitates consensus among all nodes for audit conclusions, significantly reducing the risk of malicious auditor actions and ensuring the fairness and security of copyright management. Furthermore, as participants on the blockchain, Auditors face the potential reduction in their weight following errors, which could negatively impact their standing and interests. This creates a motivation for auditors to uphold impartiality and strive for precision, ensuring that their judgments are accurate, equitable, and reliable.
Theorem 4. Parameters $m$ = 288 and $k$ = 10 are feasible and secure for CTrace under a false positive probability of 0.1%, ensuring efficient and reliable copyright verification for AIGC products.
Proof: The security and feasibility of CTrace depend on its ability to balance false positive probability $P_{fp}$, the size of array $m$, and the number of hash functions $k$. For a given $n$ (number of elements inserted), the false positive rate is bounded by:
$P_{fp}=\left(1-e^{-\frac{kn}{m}}\right)^{k}$
To achieve $P_{fp}\leq$ 0.1%, we derive the relationship between $m$, $k$, and $n$. The optimal number of hash functions $k$ minimizes $P_{fp}$ for fixed $m$ and $n$:
$k_{\text{opt}}=\frac{m}{n}\ln 2$
For $n\leq$ 20, substituting $m$ = 288 and $k$ = 10:
$P_{fp}$ $=\left(1-e^{-\frac{10\times 20}{288}}\right)^{10}$
$=\left(1-e^{-0.694}\right)^{10}$
$\approx(1-0.500)^{10}=0.500^{10}\approx 9.76\times 10^{-4}$
This yields $P_{fp}\approx$ 0.0976% , which meets the requirement $P_{fp}\leq$ 0.1%. The parameters are therefore feasible.
For $P_{fp}$ = 0.0976%, an attacker attempting to spoof $10^{4}$ AIGC products would trigger $\sim$ 9.76 false positives on average. This low rate makes systematic exploitation statistically impractical.
The bit density $\rho$ of CTrace is:
$\rho=1-e^{-\frac{kn}{m}}=1-e^{-\frac{10\times 20}{288}}\approx 0.500$
With $\rho\approx$ 50%, the filter maintains stable performance for $n\leq$20, avoiding accelerated false positive growth.
Using $k$ = 10 hash functions introduces $O(k)$ operations for insertion/query. Modern cryptographic hash functions (e.g., SHA-3) execute these operations in $\mu$s-scale, ensuring real-time compatibility with blockchain transactions.
Therefore, the parameters $m$ = 288 and $k$ = 10 satisfy $P_{fp}\leq$ 0.1% for $n\leq$ 20, maintaining security and feasibility.
Theorem 5. An attacker cannot feasibly manipulate CTrace to corrupt copyright verification proof.
Proof: The security of CTrace relies on the immutability of its bit array and the irreversibility of hash functions. To force a false positive, an attacker must either:
  • Reverse-engineer hash outputs: Cryptographic hashes (e.g., SHA-256) are preimage-resistant, making it infeasible to find an input $x^{\prime}$ that collides with a legitimate copyright entry $x$.
  • Overwhelm the filter: Saturating the filter requires inserting $n^{\prime}$ elements such that $P_{fp}$ exceeds 0.1%. For $m$ = 288, achieving $P_{fp}>$ 0.1% requires:
$n^{\prime}$ $>\frac{m\ln(1-p^{1/k})}{-k}=\frac{288\times\ln(1-(10^{-3})^{1/10})}{-10}$
$\approx\frac{288\times(-0.693)}{-10}\approx 20.0$
Since each insertion requires blockchain consensus, surreptitiously adding ( $n^{\prime}>$ 20 ) entries is detected and prevented by CTrace. Thus, the design of CTrace is intrinsically resistant to adversarial interference, ensuring the correctness of AIGC copyright verification.
7 Performance Evaluation
Within this section, we undertake an analysis and experimental validation of the performance of AIGC-Chain.
7.1 Implementation Overview
AIGC-Chain demonstrates cross-platform compatibility with mainstream blockchain platforms. To objectively evaluate system performance metrics, we constructed a private Ethereum testnet implementing a Proof-of-Authority (PoA) consensus mechanism. The blockchain core is developed in Go (Golang v1.19), while smart contracts are implemented as Solidity (v0.8.0) bytecode executed through the Ethereum Virtual Machine (EVM). Our experimental network architecture comprises 8 validator nodes and one bootnode, with each node instance containerized in isolated Docker environments (v20.10.18) to ensure resource partitioning. All network components operate on KVM-based virtual machines configured with 4GB DDR4 memory, and Ubuntu 22.04 LTS. Besides, the size of bloom filter used in CTrace was set to $m$=288, and the false positive rate was set at 0.1%.
Figure 7: Performance Analysis of AIGC-Chain
7.2 Blockchain Evaluation
The cost efficiency of AIGC-Chain is evaluated based on the amount of gas required for each contract function. We analyze computational costs for various procedures within AIGC-Chain, meticulously documenting transaction costs and execution costs. Transaction gas represents the actual amount of gas expended, whereas execution gas refers to the gas required specifically for the function execution, excluding the overhead associated with the transaction itself. In Ethereum, the gas fee consumed is specifically referred to as transaction gas, and it is only charged for functions that alter the state of the blockchain. Read-only functions, on the other hand, do not consume gas and are not subject to latency.
Fig. 7 provides a comparative analysis of the transaction gas costs associated with various procedures within the AIGC-Chain. It is observed that the gas cost for $Tx_{req}$ reaches a maximum of 217,370 gas units, which is notably higher than the costs incurred by the subsequent procedures. This increased expenditure is due to the resource-intensive write operations required for the generation of an AIGC product on the blockchain, including the registration of metadata, the generation of cryptographic hashes, and updates to the immutable ledger. Conversely, $Tx_{verification}$ incurs no transaction costs, as it is conducted off-chain through decentralized verification protocols and does not require any on-chain data writing. The metric for execution gas cost reflects the computational resource usage during the execution of smart contracts, including both algorithmic processing and on-chain data retrieval. As depicted in Figure 7, the AIGC-Chain exhibits a varied distribution of execution gas costs across its five main workflow stages. $Tx_{req}$ accounts for the highest execution costs, attributable to the intricate contract logic that involves multi-signature authentication and decentralized storage mechanisms.
Figure 8: Execution Efficiency of CTrace
7.3 Evaluation of CTrace
This research conducts a comparative evaluation of CTrace and Fast Query[9], with a focus on their respective optimization strategies for data retrieval. CTrace facilitates rapid transaction verification by leveraging a bloom filter within the world state in a systematic manner. In contrast, Fast Query optimizes search efficiency by using a single linked list to index only the copyright transactions related to product P. Traditional methods, however, necessitate exhaustive scans of the data on the blockchain for query resolution. The time complexity for executing queries using these methods is detailed in Table II, where m denotes the total number of transactions on the blockchain and n signifies the number of transactions relevant to product P. This comparative analysis quantitatively demonstrates the superiority of CTrace in data verification.
TABLE II: COMPARISON OF TIME COMPLEXITY
Method Time Complexity
Normal Query O(m)
Fast Query O(n)
CTrace O(1)
In the performance evaluation phase, we conducted rigorous comparative experiments between traditional Bloom filters and CTrace. Each experimental group involved standardized sequences of 10,000 insert and query operations separately, with each experiment repeated 30 times to eliminate systemic errors through averaging. As quantitatively demonstrated in Fig. 8, CTrace exhibits measurable throughput degradation compared to conventional Bloom filters. This performance differential primarily stems from its intrinsic privacy-preserving mechanisms. While this design decision incurs additional computational costs, it successfully prevents the avenues for retroactively inferring the original data through statistical analysis techniques. Nevertheless, this performance trade-off is deemed reasonable and acceptable within the domain of privacy-enhanced data structures.
Furthermore, we compared the performance of CTrace with traditional query methods and Fast Query, which relies on a single linked list, focusing on query time and storage overhead. Fig. 9(a) illustrates the query time costs of these three methods under varying data volumes. Traditional methods require invoking smart contracts to access transaction data on the blockchain, resulting in a consistent but significant time cost. Fast Query, with its off-chain single linked list storage, experiences increasing query times with the expansion of data. In contrast, CTrace, by only accessing the bloom filter stored in the world state and performing a single hash mapping operation, maintains a constant query time regardless of storage size. Fig. 9(b) depicts the storage costs under different data volumes, highlighting that storage requirements of CTrace remain constant and do not increase with data accumulation, in contrast to Fast Query. As the amount of data increases, CTrace will significantly reduce storage overhead compared to Fast Query, demonstrating superior space efficiency.
Figure 9: Performance Analysis of CTrace
8 Conclusion
This paper confronts the copyright challenges associated with AIGC and introduces a trustworthy AIGC copyright management system for AIGC. It meticulously records intermediate data throughout the full lifecycle of AIGC and anchors it within a decentralized blockchain system, thereby establishing a robust multi-party supervision framework for the fair and reliable management of AIGC Copyrights. The paper further develops a copyright tracing method utilizing Bloom filters, named CTrace, which markedly improves the efficiency and accuracy of blockchain transaction queries and verifications, and furnishes a solid proof base for adjudicating copyright disputes. These innovations robustly safeguard the legitimate rights and offer a reliable framework for the collection and verification of proof in copyright disputes, promoting the sustainable growth and vibrancy of AIGC. As future directions, we plan to divide the copyright management of AIGC into three distinct phases: model training, requirement formulation, and content generation. We aim to delve into the copyright concerns that may emerge at each phase and to propose effective resolution strategies.